the radius of a nucleus of mass no. 16 is 3 Fermi.Find the radius of the nucleus of mass no. 128?


 

Grade:12

3 Answers

Ramesh V
14 years ago

Average radius if nucleus is given by:

R = R0A1/3       A:mass number    R0 = 1.1x10-15m =1.1 fermi

so, R1/R2 = (A1/A2)1/3

given R1 = unknown ,    R2 = 3 fm  , A1 = 128  ,      A2 = 16

So, R1 = 2*(128/16)1/3

R1 = 6 fermi

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch

Ramesh V
14 years ago

A small mistake from above solution

its

So, R1 = 3*(128/16)1/3

R1 = 6 fermi

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Average radius if nucleus is given by:
R = R0A^(1/3) A:mass number R0 = 1.1x10-15m =1.1 fermi
so, R1/R2 = (A1/A2)^1/3
given R1 = unknown , R2 = 3 fm , A1 = 128 , A2 = 16
So, R1 = 3*(128/16)^1/3
R1 = 6 fermi

Thanks and Regards